OpenStudy (anonymous):

please help Solve for x: 2x - 2 > 4x + 6 x < -4 x > -4 x < 4 x > 4

OpenStudy (jdoe0001):

what would you get if you were to solve say 2x - 2 = 4x + 6 ? solve for "x" that is

OpenStudy (anonymous):

4

OpenStudy (jdoe0001):

well, kinda 2x - 2 = 4x + 6 # add -2x off each -2 = 2x +6 # add -6 off each -8 = 2x so x = -4 for an inequality, the factoring is exactly the same, so you'd get off that x > -4

OpenStudy (anonymous):

can you please help me more on question please

OpenStudy (anonymous):

What is the equation of the line that passes through the point (4, 1) and has a slope of 5? y = 5x - 3 y = 5x - 1 y = 5x + 3 y = 5x - 19

OpenStudy (jdoe0001):

the only thing that changes in an inequality is that when you multiply or divide or exponentialize by some negative number THEN you'd have to "flip" the sign like say if you had 8 > -2x, so 8 > -2x, # DIVIDE both sides by -2 \(\large -4 \lt x\) so, the sign would "flip" but in this case we didn't, we just add it only, but so you know :)

OpenStudy (anonymous):

okay thank you for telling me that, that was very great helful tip you gave me

OpenStudy (jdoe0001):

yw

OpenStudy (anonymous):

can u help me on other question please

OpenStudy (jdoe0001):

http://www.teachhandson.com/calculationb.gif see the example? they're using the "point-slope" form in the example they have a slope and a point on the line